OpenStudy (fluffybear):

What is the sum? –19 + 41 ________?

OpenStudy (idku):

what is 41 - 19?\

OpenStudy (idku):

You can always abide by, -a+b = b-a

OpenStudy (idku):

41 - 19 ---- ? ?

OpenStudy (fluffybear):

Is it 60

OpenStudy (idku):

41 + 19 is 60 But, 41 - 19 = what?

OpenStudy (fluffybear):

22

OpenStudy (idku):

Yes

OpenStudy (fluffybear):

but, wait is it 22 or -22?

OpenStudy (idku):

So, if 41-19=22 then -19+41=22 also

OpenStudy (idku):

Whatever number has more value, you go by.

OpenStudy (idku):

Which number has more value 41 or 19?

OpenStudy (idku):

41 is greater than 19, right?

OpenStudy (fluffybear):

Absolutely :) thank you so much for explaining

OpenStudy (idku):

So, the result gets the sign of the number with greater number (in this case, it gets + from 41).

OpenStudy (idku):

So, -19 + 41 = 22

OpenStudy (idku):

yw